Tuesday, 25 August

Special Activities

\underline{19:30 – 21:00 H 104}

International Comparison of Mathematical Studies, University Degrees, and Professional Perspectives

Round table, organized by Ina Kersten (Bielefeld, Germany), among the panelists are: D. Cioranescu, K.-H. Hoffmann, J. Netzel, A. Pelczar, M. Primicerio

ICM’98 – Program  1